OpenStudy (anonymous):

kindly help me: -4+3(x^2-2[y-3(x^2-y)]+4}

OpenStudy (angela210793):

=21x^2-24y+8 Hope i haven't forgotten anything...oO

OpenStudy (anonymous):

-4+3{x^2-2[y-3(x^2-y)]+4} =-4+3*x^2-3*2[y-3(x^2-y)]+3*4 =-4+3x^2-6[y-3(x^2-y)]+12 =-4+3x^2-6*y+18(x^2-y)+12 =-4+3x^2-6y+18x^2-18y+12 =21x^2-24y+8